40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 21 23:59:12 2018, shruti, Update, PEM, Seismometer temp control 
    Reply  Mon Jun 25 22:27:57 2018, shruti, Update, PEM, Seismometer temp control - heater circuit 
    Reply  Sun Jul 29 20:03:02 2018, rana, Update, PEM, Seismometer temp control Screen_Shot_2018-07-29_at_8.00.27_PM.png
Message ID: 14016     Entry time: Mon Jun 25 22:27:57 2018     In reply to: 14001
Author: shruti 
Type: Update 
Category: PEM 
Subject: Seismometer temp control - heater circuit 

After removing all the clamping screws from the heater circuit board, I soldered the wire connecting IRF630 to the output of OP27, which had come off earlier. This can only be a temporary fix as the wire was not long enough to be able to make a proper solder joint. I also tried fixing two other connections which were also almost breaking.

After re-assembling everything I found out that one of the LEDs was not working. The most likely cause seems to be an issue with LM791, LM 781 or the LED itself. Due to the positioning of the wires, I was unable to test them today but will try again possibly tomorrow.

Equipment used for this is still lying at the X end.

Quote:

We (Rana and I) are re-assembling the temperature controls on the seismometer to attempt PID control and then improve it using reinforcement learning.

We tried to re-assemble the connections for the heater and in-loop temperature sensor on the can that covers the seismometer.

We fixed (soldered) two of the connections from the heater circuit to the heater, but did not manage to get the PID working as one of the wires attached to the MOSFET had come off. Re-soldering the wire would be attempted tomorrow.

Equipment for undertaking all this is still left at the X-end of the interferometer and will be cleared soon.

ELOG V3.1.3-